# A neural-network based nonlinear non-intrusive reduced basis method with online adaptation for parametrized partial differential equations

### v1

@@ Abstract
We propose a nonlinear, non-intrusive reduced basis method with online adaptation for efficient approximation of parametrized partial differential equations. The method combines neural networks with reduced-order modeling and physics-informed training to enhance both accuracy and efficiency. In the offline stage, reduced basis functions are obtained via nonlinear dimension reduction, and a neural surrogate is trained to map parameters to approximate solutions. The surrogate employs a nonlinear reconstruction of the solution from the basis functions, enabling more accurate representation of complex solution structures than linear mappings. The model is further refined during the online stage using lightweight physics-informed neural network training. This offline–online framework enables accurate prediction especially in complex scenarios or with limited snapshot data. We demonstrate the performance and effectiveness of the proposed method through numerical experiments.
@@ 1 Introduction
Parametrized partial differential equations (PDEs) arise in a wide range of scientific and engineering applications, including uncertainty quantification, optimization, and control, where one must repeatedly solve the same PDE for many different parameter values [17, 28]. High-fidelity discretizations such as finite element, spectral, or finite difference methods can be prohibitively expensive in this many-query context. The reduced basis (RB) methods [17, 28] address this challenge by constructing, in an expensive offline stage, a low‐dimensional approximation space tailored to the solution manifold, and then performing online solves in this reduced space. Since the online solves are performed in a reduced space, they are often significantly more efficient than solving the full PDE directly. While the standard RB methods can yield significant speedups, they still require the solution of algebraic systems at the online stage. This step can remain computationally demanding, particularly when the underlying PDE is nonlinear.
@@
Non‐intrusive RB approaches reduce computational cost during the online stage by replacing the traditional solver with a regression model that is constructed in the offline stage and directly maps parameters to RB coefficients. While noting that early works employ conventional interpolation methods (see, e.g., [5, 33]), our focus here is on more recent developments where trained neural networks (NN) are utilized to predict RB coefficients from parameters (see, e.g., [18, 32, 10]). While efficient, these non-intrusive methods rely on a surrogate model constructed entirely during the offline stage. Consequently, if the surrogate lacks sufficient accuracy or generalizes poorly, it may result in unreliable solution predictions, with no correction mechanism available during the online stage. To address this limitation, we propose a two-stage machine learning-based RB framework, which builds upon the non-intrusive RB approach while employing an online adaptation to enhance the solution for a given parameter value. In this novel approach, a non-intrusive surrogate model is trained in the offline stage and then refined in an unsupervised manner during the online phase to improve prediction accuracy. Importantly, the online refinement process is designed to remain computationally efficient, maintaining the low-cost advantage of non-intrusive RB methods.
Another important feature of the proposed method is its use of nonlinear surrogate model. Standard RB methods and most of non-intrusive RB methods represent the solution as a linear combination of basis functions, with some exceptions such as the autoencoder-based method proposed in [12]. While linear representations are computationally convenient, they can struggle to represent strongly nonlinear solution manifolds. In contrast, the proposed method constructs an NN-based nonlinear mapping of the basis functions, which can more effectively capture complex, nonlinear solution manifolds. This mapping is learned during the offline stage and subsequently adapted in the online stage. It is also worth mentioning that our method includes two main components in the offline stage. First, the reduced basis functions are obtained via an NN-based nonlinear dimension reduction. Second, model parameters enter the formulation through a hypernetwork of the main network that maps the reduced basis functions to the approximate solution.
As discussed earlier, online training plays a central role in the proposed method. Notably, this online training is performed within the framework of physics‐informed neural networks (PINNs) [29, 3], which approximates PDE solutions by embedding residual minimization into the loss function of a neural network surrogate. A major advantage of our method is that, thanks to the offline computation, we can accurately approximate the solution with a lightweight network that can be trained efficiently. Specifically, only a small subset of the network parameters is adapted during the online stage, and the hypernetwork provides a well-informed ‘warm start’ that enables rapid convergence. More details are provided in Section 3.
To summarize, we propose an NN-based nonlinear RB method that combines the strengths of non-intrusive RB methods and online adaptation: in the offline stage, we obtain basis functions via a nonlinear dimension reduction and then train a neural network—whose weights are initialized by a parameter‐dependent hypernetwork—to nonlinearly map these basis functions to the solution; in the online stage, we refine the prediction via a lightweight online PINN‐like training. This hybrid framework provides enhanced accuracy with an efficient online training procedure.
@@
The remainder of this work is organized as follows. Section 2 presents background material, including the parametrized PDE model of interest, standard and non-intrusive RB methods, and PINNs. Section 3 introduces the proposed NN-based nonlinear RB method, outlining its offline and online components in detail. Section 4 presents numerical experiments that demonstrate the effectiveness of the proposed method. Finally, Section 5 concludes the paper with a summary and potential directions for future research.
2 Background
@@ 2.1 Problem Setup
${\mathcal{N}}\bigl(u(\mathbf{x};\boldsymbol{\mu});\boldsymbol{\mu}\bigr)=0,\quad\mathbf{x}\in\Omega,$ (2.1a)
subject to a boundary condition (BC):
${\mathcal{B}}\bigl(u(\mathbf{x};\boldsymbol{\mu});\boldsymbol{\mu}\bigr)=0,\quad\mathbf{x}\in\partial\Omega.$ (2.1b)
Here, $\mathcal{N}$} and $\mathcal{B}$ are two differential operators, $u(\mathbf{x};\boldsymbol{\mu})$ is the solution, $\mathbf{x}$ is the $d$-dimensional spatial variable, and $\Omega\subset\mathbb{R}^{d}$ is the spatial domain with $\partial\Omega$ being its boundary; $\boldsymbol{\mu}\in\mathcal{Y}$ represents the collection of model parameters, with $\mathcal{Y}$ being the parameter domain. For the sake of convenience, throughout this paper, we assume that the system consists of one primary PDE and one BC equation; however, the methods and techniques discussed here are applicable to systems with multiple PDEs and BC equations.
In many practical settings, $\boldsymbol{\mu}$ is modelled as a random variable with a prescribed probability distribution $\pi_{\boldsymbol{\mu}}(\cdot)$. Our objective is to numerically solve the parametrized PDE for a given value of $\boldsymbol{\mu}$. Conventionally, such PDEs are solved using standard numerical schemes such as the finite difference method (FDM), the finite element method (FEM), or spectral methods. Here, we present the Galerkin projection based FEM as an example (see, e.g., [19]) . Let $\mathcal{V}$ be a Hilbert space equipped with the inner product $(\cdot,\cdot)_{\mathcal{V}}$ and norm $\|\cdot\|_{\mathcal{V}}$. Its dual space is denoted by $\mathcal{V}^{\prime}$, and the duality pairing is given by $\langle\cdot,\cdot\rangle_{\mathcal{V}^{\prime},\mathcal{V}}:\mathcal{V}^{\prime}\times\mathcal{V}\rightarrow\mathbb{R}$. We can view the differential operator $\mathcal{N}$} in (2.1) as a mapping:
${\mathcal{N}}:\mathcal{V}\times\mathcal{Y}\longrightarrow\mathcal{V}^{\prime}.$ (2.2)
For a fixed parameter $\overline{\boldsymbol{\mu}}\in\mathcal{Y}$, the weak formulation of (2.1) is: find $u(\cdot;\overline{\boldsymbol{\mu}})\in\mathcal{V}$ such that
$\langle{\tilde{\mathcal{N}}}\bigl(u(\cdot;\overline{\boldsymbol{\mu}});\overline{\boldsymbol{\mu}}\bigr),v\rangle_{\mathcal{V}^{\prime},\mathcal{V}}=0\quad\forall v\in\mathcal{V},$ (2.3)
with $\tilde{\mathcal{N}}$ representing both the operator $\mathcal{N}$ in Eq. (2.1) and the BC operator $\mathcal{B}$ in Eq. (2.1b).
Next the solution $u\in\mathcal{V}$ is approximated in a finite-dimensional subspace $\mathcal{V}_{n}\subset\mathcal{V}$, spanned by $n$ basis functions $\{\eta_{1},\eta_{2},\ldots,\eta_{n}\}$. The approximate solution is expressed as a linear combination:
$u_{h}(\boldsymbol{x};\overline{\boldsymbol{\mu}})=\sum_{j=1}^{n}\tilde{u}_{j}(\overline{\boldsymbol{\mu}})\,\eta_{j}(\boldsymbol{x}),$ (2.4)
@@
$\tilde{\mathbf{u}}(\overline{\boldsymbol{\mu}})=\begin{bmatrix}\tilde{u}_{1}(\overline{\boldsymbol{\mu}})&\ldots&\tilde{u}_{n}(\overline{\boldsymbol{\mu}})\end{bmatrix}^{T}\in\mathbb{R}^{n},$ (2.5)
@@
$\langle\tilde{\mathcal{N}}\bigl(u_{h}(\cdot;\overline{\boldsymbol{\mu}});\overline{\boldsymbol{\mu}}\bigr),v_{h}\rangle_{\mathcal{V}^{\prime},\mathcal{V}}=0\quad\forall v_{h}\in\mathcal{V}_{n}.$ (2.6)
Since $\mathcal{V}_{n}$ is spanned by basis functions $\{\eta_{1},\eta_{2},\ldots,\eta_{n}\}$, Eq. (2.6) is equivalent to
$\langle\tilde{\mathcal{N}}\bigl(u_{h}(\cdot;\overline{\boldsymbol{\mu}});\overline{\boldsymbol{\mu}}\bigr),\eta_{j}\rangle_{\mathcal{V}^{\prime},\mathcal{V}}=0,\,j=1,\ldots,n.$ (2.7)
Next we need to solve the equation system (2.7). If $\tilde{\mathcal{N}}$ is a linear operator, it reduces to an $n\times n$ linear system. If $\tilde{\mathcal{N}}$ is nonlinear, a common approach is to solve the equation system using a linearization procedure (e.g., the Newton’s method [11]).
@@
2.2 Reduced basis method
In this section, we give a brief overview of the RB method. The main idea of the method is rather straightforward: first, we construct a low-dimensional approximation space that is specifically tailored to capture the essential features of the solution manifold across the parameter space, and then we solve the PDE for a specific parameter value in this reduced low-dimensional space, significantly lowering computational costs compared to the standard approaches. The standard RB method is implemented in two stages: an offline stage and an online stage. In the offline stage, the reduced-dimensional space for solution approximation is constructed. This stage is computationally expensive but performed only once. In the online stage, the reduced space obtained offline is used to rapidly approximate solutions for new parameter values, enabling real-time simulations of the PDE model.
@@ 2.2.1 Offline stage
${u}_{h}(\cdot;\boldsymbol{\mu}^{(i)}),\quad i=1,2,\dots,n_{s};$ (2.8)
@@
where $\mathbf{B}\in\mathbb{R}^{n\times n}$, $\mathbf{\Sigma}\in\mathbb{R}^{n\times n_{s}}$ (diagonal matrix containing singular values), and $\mathbf{V}\in\mathbb{R}^{n_{s}\times n_{s}}$. We select the first $r\ll n$ singular vectors corresponding to the largest singular values to construct the reduced basis functions:
@@
where $\{\eta_{j}(\mathbf{x})\}_{j=1}^{n}$ are the original basis functions from the high-fidelity solution space. The reduced basis space is thus defined by
$\mathcal{V}_{r}=\mathrm{span}\{\psi_{1},\psi_{2},\dots,\psi_{r}\},$ (2.12)
and in the online stage, that is described next, we will compute the reduced solution as an element of $\mathcal{V}_{r}$ .
@@ 2.2.2 Online stage
Once the reduced space $\mathcal{V}_{r}=\text{span}\{\psi_{1},\dots,\psi_{r}\}$ has been constructed in the offline stage, the online stage aims to compute an efficient approximation of the high-fidelity solution for a new parameter value ${\boldsymbol{\mu}}^{*}\in\mathcal{Y}$. We seek a reduced solution $u_{\rm rb}(\mathbf{x},\boldsymbol{\mu}^{*})\in\mathcal{V}_{r}$ expressed as a linear combination of the reduced basis functions:
@@
$\langle\tilde{\mathcal{N}}\bigl(u_{\rm rb}(\cdot,{\boldsymbol{\mu}}^{*});{\boldsymbol{\mu}}^{*}\bigr),\psi_{j}\rangle_{\mathcal{V}^{\prime},\mathcal{V}}=0,\quad j=1,\ldots,r,$ (2.15)

### v2

# arXiv:2511.07684v2[math.NA] 02 Mar 2026

@@ Abstract
We introduce a Nemytskii neural operator framework for nonlinear model reduction of parametrized steady-state partial differential equations. The method generalizes reduced basis approaches by replacing linear combinations of basis functions with a structured nonlinear mapping realized through a pointwise Nemytskii operator acting on fixed feature functions. Feature functions are learned offline via nonlinear dimension reduction from high-fidelity snapshots, and a hypernetwork maps model parameters to a lightweight reconstruction network, which is further refined online using physics-informed residual minimization. The Nemytskii structure preserves analytical regularity and enables efficient evaluation of spatial and parametric derivatives, leading to fast online adaptation. Numerical experiments demonstrate that the proposed method consistently outperforms linear model reduction techniques, particularly for complex solution manifolds.
@@ 1 Introduction
Parametrized partial differential equations (PDEs) arise in a wide range of scientific and engineering applications, including uncertainty quantification, optimization, and control, where one must repeatedly solve the same PDE for many different parameter values [16, 28]. High-fidelity discretizations such as finite element, spectral, or finite difference methods can be prohibitively expensive in this many-query context. The reduced basis (RB) methods [16, 28] address this challenge by assuming that the solution manifold can be well approximated by a low-dimensional linear subspace of the solution space. In an expensive offline stage, RB methods construct a reduced approximation space (typically from high-fidelity snapshots); in the online stage, they compute solutions by solving the governing equations projected onto this reduced space. Because the online solves involve only a small number of degrees of freedom, they are often significantly more efficient than solving the full PDE directly.
@@
While the standard RB methods can yield significant speedups, they still require the solution of algebraic systems at the online stage. This step can remain computationally demanding, particularly when the underlying PDE is nonlinear. Non‐intrusive RB approaches reduce computational cost during the online stage by replacing the traditional solver with a regression model that is constructed in the offline stage and directly maps parameters to RB coefficients. While noting that early works employ conventional interpolation methods (see, e.g., [4, 33]), our focus here is on more recent developments where trained neural networks (NN) are utilized to predict RB coefficients from parameters (see, e.g., [17, 32, 9]).
It is important to emphasize that all RB-type methods rely on the central assumption that the solution manifold admits a good approximation by a low-dimensional linear subspace; in practice, this assumption may not hold, or the dimension of the required linear subspace may be unacceptably high. When linear approximation is not effective, nonlinear reduced-order modeling (ROM) provides a natural alternative framework. These methods (e.g., [20, 11]) seek to represent the solution set using nonlinear low-dimensional structures, which can capture complex solution behavior more efficiently than linear subspaces. In contrast to RB methods, which construct reduced spaces from explicitly defined basis functions and thus impose a clear mathematical structure, many existing nonlinear ROM approaches rely on fully data-driven models in which the solution representation and its structure are learned primarily from data. As a result, such methods often require large training datasets and complex neural networks, leading to increased training cost and limited robustness and generalization. A more detailed discussion of nonlinear ROM methods and their limitations is provided in Section 2.4.
Motivated by these considerations, we introduce a Nemytskii neural operator (NNO) framework for nonlinear ROM of parametrized PDEs. The proposed approach can be regarded as a nonlinear generalization of RB methods: instead of representing the solution as a linear combination of basis functions, it constructs the approximation through a nonlinear mapping realized by a pointwise Nemytskii operator acting on fixed feature functions. The pointwise (superposition) nature of the Nemytskii operator ensures that the solution is constructed locally from the fixed feature functions; as a result, the resulting nonlinear mapping is highly structured and principled, with its structure strongly constrained by the feature functions. This property is closely analogous to the role played by basis functions in RB methods. As such the proposed NNO framework overcomes the limitations of linear approximation in RB methods, while retaining their ability to impose explicit analytical structure through fixed feature functions, a key distinction from generic nonlinear ROM.
Beyond the Nemytskii representation, the proposed NNO framework consists of two additional key components that are essential to its performance: online adaptation and NNO-based dimension reduction.
Many learning–based ROM/RB methods, such as non-intrusive RB approaches, rely on surrogate models that are constructed entirely during the offline stage. In the online stage, the parameter value of interest is simply provided as input to the trained model to obtain an approximate solution. The surrogate models constructed offline may not be sufficiently accurate, particularly in parameter regimes that are poorly represented in the training data, leading to unreliable solution predictions with no correction mechanism available during the online stage. To address this issue, the proposed NNO framework incorporates an online adaptation mechanism that allows the solution approximation to be adapted for each new parameter instance during the online stage. The adaptation is performed in an unsupervised manner using physics-based residuals associated with the governing PDE [29, 2]. Importantly the NNO framework allows us to represent the solution using a lightweight network defined on the feature space, enabling efficient adaptation during the online stage.
Another important component of the proposed method is nonlinear dimension reduction (feature extraction) performed at the offline stage. The purpose of this step is to learn a set of feature functions from high-fidelity solution snapshots. In the proposed method, the feature functions are also constructed through a Nemytskii-operator–based nonlinear mapping applied to the snapshot data. This design enables the model to effectively learn the underlying analytical structure from the solution snapshots. The learned feature functions are fixed during the online stage to ensure computational efficiency.
In summary, the main contribution of this paper is the introduction of a Nemytskii neural operator (NNO) framework for constructing a nonlinear approximation of the solution manifold of parametrized PDEs. The proposed approach combines structured nonlinear solution representation with offline nonlinear dimension reduction and efficient online adaptation, enabling expressive approximation while retaining analytical structure and computational efficiency. By exploiting the pointwise nature of the Nemytskii operator and operating in a low-dimensional feature space, the method provides a principled alternative to both linear RB methods and generic data-driven nonlinear ROM approaches.
@@
The remainder of this work is organized as follows. Section 2 presents background and preliminary material, including the parametrized PDE model of interest, standard and non-intrusive RB methods, nonlinear ROM and physics-informed neural networks (PINNs). Section 3 introduces the proposed NNO method, outlining its offline and online components in detail. Section 4 presents numerical experiments that demonstrate the effectiveness of the proposed method. Finally, Section 5 concludes the paper with a summary and potential directions for future research.
2 Preliminaries and background
@@ 2.1 Problem Setup
${\mathcal{N}}\bigl(u(\mathbf{x},\boldsymbol{\mu}),\boldsymbol{\mu}\bigr)=0,\quad\mathbf{x}\in\Omega,$ (2.1a)
subject to Dirichlet (essential) boundary condition (BC):
$u(\mathbf{x},\boldsymbol{\mu})=b(\mathbf{x},\boldsymbol{\mu})\quad\mathbf{x}\in\partial\Omega.$ (2.1b)
Here, $\mathcal{N}$} is a (possibly nonlinear) differential operator, $u(\mathbf{x},\boldsymbol{\mu})$ is the solution, $\mathbf{x}$ is the $d$-dimensional spatial variable, and $\Omega\subset\mathbb{R}^{d}$ is the spatial domain with $\partial\Omega$ being its boundary; $\boldsymbol{\mu}\in{Y}$ represents the collection of model parameters, with Y{Y} being the parameter domain. Since natural boundary conditions can be incorporated implicitly in the weak formulation and therefore do not need to be treated explicitly in the methods considered in this work, we restrict our attention to Dirichlet BCs.
In many practical settings, $\boldsymbol{\mu}$ is modelled as a random variable with a prescribed probability distribution $\pi_{\boldsymbol{\mu}}(\cdot)$. Our objective is to numerically solve the parametrized PDE for a given value of $\boldsymbol{\mu}$. Conventionally, such PDEs are solved using standard numerical schemes such as the finite difference method (FDM), the finite element method (FEM), or spectral methods. Here, we present the Galerkin FEM as an example (see, e.g., [18]) . Let $\mathcal{V}$ be a Hilbert space equipped with the inner product $\langle\cdot,\cdot\rangle_{\mathcal{V}}$ and norm $\|\cdot\|_{\mathcal{V}}$. We assume that $\mathcal{V}$ is chosen to satisfy the Dirichlet boundary conditions. Its dual space is denoted by $\mathcal{V}^{\prime}$, and the duality pairing is given by $\langle\cdot,\cdot\rangle_{\mathcal{V}^{\prime},\mathcal{V}}:\mathcal{V}^{\prime}\times\mathcal{V}\rightarrow\mathbb{R}$. We can view the differential operator $\mathcal{N}$} in (2.1) as a mapping:
${\mathcal{N}}:\mathcal{V}\times{Y}\longrightarrow\mathcal{V}^{\prime}.$
The weak formulation of (2.1) is:
$\langle{{\mathcal{N}}}\bigl(u(\cdot,{\boldsymbol{\mu}}),{\boldsymbol{\mu}}\bigr),v\rangle_{\mathcal{V}^{\prime},\mathcal{V}}=0,\quad\forall v\in\mathcal{V},$ (2.2)
and a solution to Eq. (2.2) is called a weak solution to Eq. (2.1).
Next the weak solution $u\in\mathcal{V}$ is constructed in a finite-dimensional subspace $\mathcal{V}_{n}\subset\mathcal{V}$, spanned by $n$ basis functions $\{\eta_{1},\eta_{2},\ldots,\eta_{n}\}$. The numerical solution is expressed as a linear combination:
$u_{h}(\mathbf{x},{\boldsymbol{\mu}})=\sum_{j=1}^{n}\tilde{u}_{j}({\boldsymbol{\mu}})\,\eta_{j}(\mathbf{x}),$ (2.3)
@@
$\tilde{\mathbf{u}}({\boldsymbol{\mu}})=\begin{bmatrix}\tilde{u}_{1}({\boldsymbol{\mu}})&\ldots&\tilde{u}_{n}({\boldsymbol{\mu}})\end{bmatrix}^{T}\in\mathbb{R}^{n},$ (2.4)
@@
$\langle{\mathcal{N}}\bigl(u_{h}(\cdot,{\boldsymbol{\mu}}),{\boldsymbol{\mu}}\bigr),\eta_{j}\rangle_{\mathcal{V}^{\prime},\mathcal{V}}=0,\,j=1,\ldots,n.$ (2.5)
Next we need to solve the equation system (2.5). If $\mathcal{N}$} is a linear operator, it reduces to a linear system. If $\mathcal{N}$} is nonlinear, a common approach is to solve the equation system using a linearization procedure (e.g., Newton’s method [10]).
@@
2.2 Standard reduced basis method
In this section, we give a brief overview of the RB method. The main idea of the method is rather straightforward: first, we construct a low-dimensional approximation space that can well approximate the solution set across the parameter space, and then we solve the PDE for a specific parameter value in this reduced low-dimensional space, significantly lowering computational costs compared to the standard approaches. The standard RB method is implemented in two stages: an offline stage and an online stage. In the offline stage, a reduced-dimensional space for solution approximation is constructed. This stage is computationally expensive but performed only once. In the online stage, the reduced space obtained offline is used to rapidly approximate solutions for new parameter values, enabling real-time simulations of the PDE model.
2.2.1 Low-dimensional linear subspace
Let the solution set of Eq. (2.2) be,
$\mathcal{M}:=\{\,u(\cdot,\boldsymbol{\mu})\in\mathcal{V}\mid\boldsymbol{\mu}\in Y\,\}\subset\mathcal{V},$
and we assume that $\mathcal{M}$ can be well approximated by a $r$-dimensional linear subspace:
$\mathcal{V}_{r}=\mathrm{span}\{\psi_{1},\psi_{2},\dots,\psi_{r}\},$ (2.6)
where $\psi_{1},...,\psi_{r}$ are the basis functions spanning $\mathcal{V}_{r}$. That is, each $u(\cdot,\boldsymbol{\mu})\in\mathcal{M}$ can be well approximated by a linear combination of $\psi_{1},...,\psi_{r}$. For RB methods to be effective, the reduced dimension $r$ needs to be much smaller than $n$, the dimensionality of the FEM space. A key step in RB methods is to obtain the basis functions $\psi_{1},...,\psi_{r}$, which are constructed in the offline stage.
@@ 2.2.2 Offline stage
${u}_{h}(\cdot,\boldsymbol{\mu}^{(i)}),\quad i=1,2,\dots,N_{s};$ (2.7)
@@
As an alternative to random sampling, greedy algorithms are also often employed to adaptively generate snapshots; see, e.g., [28, 16].
@@
where $\mathbf{B}\in\mathbb{R}^{n\times n}$, $\mathbf{\Sigma}\in\mathbb{R}^{n\times N_{s}}$ (diagonal matrix containing singular values), and $\mathbf{V}\in\mathbb{R}^{N_{s}\times N_{s}}$. We select the first $r\ll n$ singular vectors corresponding to the largest singular values to construct the basis functions of $\mathcal{V}_{r}$:
@@
where $\{\eta_{j}(\mathbf{x})\}_{j=1}^{n}$ are the original basis functions from the high-fidelity solution space.
@@ 2.2.3 Online stage
Once the reduced space $\mathcal{V}_{r}=\text{span}\{\psi_{1},\dots,\psi_{r}\}$ has been constructed in the offline stage, the online stage aims to compute an efficient approximation of the high-fidelity solution for a new parameter value ${\boldsymbol{\mu}}^{*}\in{Y}$. We seek a reduced solution $u_{\rm rb}(\mathbf{x},\boldsymbol{\mu}^{*})\in\mathcal{V}_{r}$ expressed as a linear combination of the reduced basis functions:
@@
$\langle{\mathcal{N}}\bigl(u_{\rm rb}(\cdot,{\boldsymbol{\mu}}^{*}),{\boldsymbol{\mu}}^{*}\bigr),\psi_{j}\rangle_{\mathcal{V}^{\prime},\mathcal{V}}=0,\quad j=1,\ldots,r,$ (2.13)